OpenStudy (anonymous):

The expression cosx + tanx sinx can be simplified to _____. A. cosx sinx B. secx C. cotx D. cscx

OpenStudy (tkhunny):

tan(x) = sin(x)/cos(x) Common Denominator. Numerator should be interesting.

OpenStudy (anonymous):

is it B

OpenStudy (tkhunny):

You tell me. How did you conclude that?

OpenStudy (anonymous):

yes, B.\[cosx +tanxsinx= cosx + \frac{ sinx }{ cosx }(sinx)=cosx + \frac{ \sin^2x }{ cosx }=\frac{ \cos^2x+\sin^x }{ cosx }=\frac{ 1 }{ cosx }=secx\]

OpenStudy (tkhunny):

{sigh}